# Conditioned quantum-assisted deep generative surrogate for particle-calorimeter interactions

### v1

@@ I Introduction
By the end of the decade, the LHC is expected to begin an upgraded “High Luminosity” phase, which will ultimately increase the collision rate by a factor of 10 higher than the initial design. Increasing the number of collisions will generate more experimental data, enabling the observation of rare processes and increased precision in measurements, furthering our understanding of the universe. The path toward the HL-LHC presents great technological challenges and commensurate innovations to overcome them. Monte Carlo simulations of collision events at the ATLAS experiment have played a key role in the design of future experiments and, particularly, in the analysis of current ones. However, these simulations are computationally intensive, projected to reach into millions of CPU-years per year during the HL-LHC run [1]. Simulating a single event with Geant4 [4] in an LHC experiment requires roughly 1000 CPU seconds. The calorimeter simulation is by far dominating the total simulation time [2]. To address this challenge, deep generative models are being developed to act as particle-calorimeter interaction surrogates, with the potential to reduce the simulation overall time by orders of magnitude. The key point to take into consideration is that one particle impacting a calorimeter can lead to thousands of secondary particles, collectively known as showers, to be tracked through the detector, while only the total energy deposit per sensitive element (a cell) is actually measured in the experiment. Hence, through the generation of these showers, non-negligible computational resources are being employed in the detailed recording of the path of these particles. The problem being addressed is whether one can bypass the path-tracing step in the simulation and generate the cell energy deposits directly from a set of well-defined parameters (e.g., type of particle, incidence energy, incidence angle, etc.) via sampling from a deep generative model.
@@
There is a large and growing body of literature addressing this critical problem via deep generative models. The earliest methods developed were Generative Adversarial Networks [5, 6, 7], which are now an integral part of the simulation pipeline [8, 9] of some experiments. Different deep generative frameworks have been proposed since then, including VAEs [10, 10, 11, 12], Normalizing Flows [13, 14], Transformers [15], Diffusion models [16, 17, 18] and combinations thereof [19, 20]. A noteworthy development in the field is the CaloChallenge-2022 endeavour, which not only catalyzed research efforts but also has enabled a better quantitative comparison across different frameworks. Defining benchmarks and metrics has also been a rather active topic of research [21, 22]. A common feature of these models is the fast generation of showers via sampling on GPUs, but futher speed increases may still be possible with alternative computing paradigms. In our work, we develop deep generative models which can be naturally encoded onto Quantum Annealers (QAs), allowing for potentially significant improvements in the speed of shower simulation.
@@ II.1 Data preprocessing
Before feeding the shower and incident energy data to the model, we apply several transformations to the shower and the incident particle energy on-the-fly. Given an event shower, $\bm{v}$, and corresponding incident energy, $e$, we first reduce the voxel energy, $v_{i}$, per event by dividing it by the incident energy, $e$, viz. $E_{i}=v_{i}/e$. Notice that $E_{i}\in\left[0,1\right]$, where the left and right bounds correspond to when the voxel energy is zero and equal to the incident energy, respectively. To remove the strict bounds, we define $u_{i}=\delta+(1-2\delta)E_{i}$, where $\delta=10^{-7}$, to prevent discontinuities during the logit transformation Specifically, we use the transformation $x_{i}=\ln u_{i}/(1-u_{i})-\ln\delta/(1-\delta)$, where the second term preserves the zero values in the transformed variable, i.e., when the voxel energy is zero, $v_{i}=0$, the transformed variable $x_{i}=0$.
@@
Besides the reconstruction of the event, the encoded data is also used to train the 4-partite RBM via the Kullback-Liebler (KL) divergence. As in traditional VAEs (full derivations can be found in Ref. [34] and in Appendix A), we optimize the evidence lower bound (ELBO) to train the Calo4pQVAE. Explicitly, the ELBO function is:
@@
The first term in Eq. (2) represents the reconstruction accuracy and, the second term corresponds to the KL divergence. To define a functional form for $p_{\theta}(\hat{\bm{x}}|\bm{z})$, certain assumptions about the distribution are typically made. A common assumption is that the likelihood of the reconstruction is Gaussian distributed, which simplifies to optimizing the mean squared error (MSE).
@@
Maximizing the log-likelihood of $p_{\theta}(\bm{\chi},\bm{\xi}|\bm{z})$ is rather difficult due to the factor containing the Dirac delta function. It may well be possible to replace the Dirac delta function with a smooth parameterized function. Yet, there is a simpler way where we instead mask $\bm{\chi}$ with $\bm{\xi}$, in which case we can neglect the term containing the Dirac delta function. The key point to stress here is that by means of the mask we split the tasks between generating zeroes and non-zeroes, and generating the voxel energy via $\bm{\chi}$ and $\bm{\xi}$, respectively. Therefore, the joint distribution in Eq. (3) becomes  [^2] We removed the $\xi$ coefficient due to normalization.:
@@
Let us focus on the first term on the r.h.s. of Eq. (7). When taking the gradient of the entropy with respect to the model parameters one faces two issues: i) due to the discrete nature of the latent variables $z$, the gradient becomes singular; and ii) when attempting to take the gradient over a discrete estimator, it is ill-defined how to take the gradient of $\sum_{\bm{z}\sim q_{\phi}(\bm{z}|\bm{x})}(...)$ with respect to $\phi$, since the gradient variables appear in the summation argument. The first issue can easily be addressed by simply considering a continuous step-like function, such as a sigmoid, $\sigma(\bullet)$. Specifically, we replace $\bm{z}$ with $\bm{\zeta}$, where $\zeta_{i}=\sigma(\left[\text{encoded data point}\right]_{i}\cdot\beta)$, and $\beta$ is an annealing parameter, such that $\lim_{\beta\rightarrow\infty}\bm{\zeta}=\bm{z}$. The second issue can be addressed by means of the so-called Gumbel trick [36]. The Gumbel trick has become an umbrella term which refers to a set of methods to sample from discrete probabilities or to estimate its partition function. In our case, we simply generate latent variables $\bm{\zeta}$ via
@@
The first term in the r.h.s. is the energy function averaged over the approximate posterior, whereas the second term is the free energy, which is independent of the encoded samples. The free energy can be replaced by the internal energy, $U$, minus the entropy, $S$, times the temperature, $T=1$, viz. $U-TS$. We use the high temperature gradient approximation to replace the second term in the r.h.s. of Eq. (10) with the internal energy. In the high temperature gradient approximation, as the energy in the system saturates, the specific heat converges to zero. We demostrate in Appendix D that when $\langle E(\bm{z})\rangle\langle\partial E(\bm{z})/\partial\phi\rangle=\langle E(\bm{z})\partial E(\bm{z})/\partial\phi\rangle$, the specific heat is zero, which allows us to rewrite the previous Equation as
@@
We then propose a Gibbs sampling process to sample from the joint distribution $p(\mathbf{v},\mathbf{h},\mathbf{s},\mathbf{t})$ of the 4-partite RBM. Notice that the joint distribution can be factorized as $p(\mathbf{v},\mathbf{h},\mathbf{s},\mathbf{t})=p(\mathbf{v}|\mathbf{h},\mathbf{s},\mathbf{t})p(\mathbf{h},\mathbf{s},\mathbf{t})$. Extending the approach used in 2-partite RBM, we perform Markov chain Monte Carlo by sequentially updating each partition conditioned on the others, i.e., the prior in step $n+1$ is be the posterior in step $n$. The Gibbs sampling process, at iteration $n$, is done in the following steps:
@@
Notice that one block Gibbs sampling step corresponds to 4 sampling steps. By repeating this process iteratively, we generate samples that approximate the joint distribution.
We have shown how to generalize an RBM to a four-partite graph. It is important to emphasize that, as with any deep generative model, the features of the dataset are expressed in latent space. In this context, each state in latent space encodes specific features of the dataset. Different techniques such as t-SNE [39] and LSD [40], can illustrate how dataset features are embedded in latent space after training. These methods allow for the conditioning of the prior on specific features. There is a rather straightforward way to condition the prior on given dataset parameters during training, which is widely used across different frameworks. Generally, this involves using specific channels to condition the prior and specifics depend on the actual framework. An accurate and robust method to sample from latent space constrained to specific features is crucial for practical applications in generative deep models. In the present context, the calorimeter surrogate requires input related to particle type, incident particle energy and incident angle, among other parameters. In the following, we show how to condition the 4-partite RBM sampling process.
@@ II.3.1 Conditioned 4-partite Restricted Boltzmann Machine
The block Gibbs sampling procedure is the same as before via Eqs. (17) although we skip Eq. (17a). The authors in [41] caution on the use of contrastive divergence when training a CRBM due to the possibility of vanishing gradient even when the conditioned RBM might not be fully trained. This happens when mixing times are larger the the number of Gibbs sampling steps. Despite the previous, we use contrastive divergence in our framework and we show how the estimated log-likelihood saturates after $150$ epochs. We further elaborate and discuss this approach in the discussion section.
@@ II.4 Quantum Annealers
A quantum annealer (QA) is an array of superconducting flux quantum bits with programmable spin–spin couplings and biases with an annealing parameter [42]. The motivation for QAs comes from the adiabatic approximation [43], which asserts that if a quantum system is in an eigenstate of its Hamiltonian (which describes the total energy of the system), and the Hamiltonian changes slowly enough, then the system will remain in an eigenstate of the Hamiltonian, although the state itself may change (see Appendix F.1 for a formal derivation). QAswere initially thought of as a faster method to find the ground state of complex problems that could be mapped onto a Hamiltonian $H$ [44]. This can be done by initializing the system in the ground state of some Hamiltonian $H_{0}$, which is easy to prepare both theoretically and experimentally. In addition, by design the commutator $[H,H_{0}]\neq 0$. The QA interpolates between the two Hamiltonians via
@@
Notice that the RBM data are binary vectors such that each element can take value of 1, whereas the qubits can have values $-1$ or $1$. Hence, to map the RBM onto the QA one redefines the RBM variables $x_{i}\rightarrow(\sigma_{z}^{(i)}+1)/2$ and rearrange the terms in the RBM Hamiltonian to be mapped onto the QA (this is explicitly shown in Appendix F). The previous variable change will lead to an energy offset between the RBM and the QA, which henceforth we neglect. It is important to stress that the QA is in a heat bath with temperature $T_{QA}\lesssim 15mK$ [52], while the annealing parameter has an upper bound $\mathcal{B}(1)\approx 5.0\cdot 10^{-24}J$ [53], which implies that the prefactor $\beta_{QA}=\frac{\mathcal{B}(1)}{2k_{B}T_{QA}}\approx 12$ while in the RBM by design $\beta_{RBM}=1$. To ensure that both, the RBM and the QA describe the same Boltzmann distribution with the same temperature, one can either re-scale the QA Hamiltonian by $1/\beta_{QA}$ or the RBM Hamiltonian by $\beta_{QA}$. Currently, there is not a way to measure the prefactor, however there are different ways to estimate it [46, 54]. In addition, there is not a direct way to control this parameter. However, the most common way to estimate this factor is via the Kullback-Liebler divergence between the QA and the RBM distributions, whereby one introduces a tuning parameter $\beta$ as a prefactor in the RBM distribution which is tuned to minimize the Kullback-Liebler divergence. It is in the process of estimating the $\beta_{QA}$ that we also indirectly control the parameter by rescaling the Hamiltonian. The previous can be mathematically expressed as
@@
The previous converges when the average energy from the RBM matches that from the QA, in which case, $\beta=\beta_{QA}$. Due to coupling-dependent temperature fluctuations, $\beta_{QA}$ is expected to change in the process of training the model. For this reason the method described by Eq. (22) should be employed, in principle, after each parameter update. Notice that optimizing $\beta$ in Eq. (22) requires samples from the RBM, furthermore, it is the classical RBM temperature which is being tuned to match the distribution of the QA. This can lead to some undesired situations, for instance, given a trained classical RBM, one might desire to fit the QA distribution to that of the RBM and not the other way around. To address this point, $H(x)$ is replaced with an annealed Hamiltonian $H(x,\beta)=H(x)/\beta$, i.e., instead of tuning the temperature in the classical RBM, to match the distribution in the QA, one iteratively rescales the Hamiltonian in the QA to effectively tune the QA’s temperature to match that of the classical RBM. This approach, as already mentioned, is well-known [46, 54, 30] and has been proven to be empirically robust yet slow to converge. We therefore propose a different protocol which empirically converges faster than the KL divergence:
@@
Up to this point, we have detailed the process of replacing the RBM with the QA in the Calo4pQVAE. As previously mentioned, we utilize a conditionalized RBM to enable the sampling of showers with specific characteristics. In the next section, we will explore various approaches for conditionalizing the QA, aiming to achieve targeted sampling of particle showers with desired features.

### v2

@@ I Introduction
By the end of the decade, the LHC is expected to begin an upgraded “High Luminosity” phase, which will ultimately increase the collision rate by a factor of 10 higher than the initial design. Increasing the number of collisions will generate more experimental data, enabling the observation of rare processes and increased precision in measurements, furthering our understanding of the universe. The path toward the HL-LHC presents great technological challenges and commensurate innovations to overcome them. Monte Carlo simulations of collision events at the ATLAS experiment have played a key role in the design of future experiments and, particularly, in the analysis of current ones. However, these simulations are computationally intensive, projected to reach into millions of CPU-years per year during the HL-LHC run [1]. Simulating a single event with Geant4 [4] in an LHC experiment requires roughly 1000 CPU seconds. The calorimeter simulation is by far dominating the total simulation time [2]. To address this challenge, deep generative models are being developed to act as particle-calorimeter interaction surrogates, with the potential to reduce the simulation overall time by orders of magnitude. The key point to take into consideration is that one particle impacting a calorimeter can lead to thousands of secondary particles, collectively known as showers, to be traced through the detector, while only the total energy deposit per sensitive element (a cell) is actually measured in the experiment. Hence, through the generation of these showers, non-negligible computational resources are being employed in the detailed recording of the path of these particles. The problem being addressed is whether one can bypass the path-tracing step in the simulation and generate the cell energy deposits directly from a set of well-defined parameters (e.g., type of particle, incidence energy, incidence angle, etc.) via sampling from a deep generative model.
@@
There is a large and growing body of literature addressing this critical problem via deep generative models. The earliest methods developed to address this challenge were of the kind of Generative Adversarial Networks [5, 6, 7], which are now an integral part of the simulation pipeline [8, 9] of some experiments. Different deep generative frameworks have been proposed since then, including VAEs [10, 10, 11, 12], Normalizing Flows [13, 14], Transformers [15], Diffusion models [16, 17, 18] and combinations thereof [19, 20]. A noteworthy development in the field is the CaloChallenge-2022 endeavour [21], which not only catalyzed research efforts but has also enabled better quantitative comparison across different frameworks. Similarly, defining benchmarks and metrics has been a rather active topic of research [22, 23]. A common feature of these models is the fast generation of showers via sampling on GPUs, but further speed increases may still be possible with alternative computing paradigms. In our work, we develop deep generative models which can be naturally encoded onto Quantum Annealers (QAs), allowing for potentially significant improvements in the speed of shower simulation.
@@ II.1 Data preprocessing
Before feeding the shower and incident energy data to the model, we apply several transformations to the shower and the incident particle energy on-the-fly. Given an event shower, $\bm{v}$, and corresponding incident energy, $e$, we first reduce the voxel energy, $v_{i}$, per event by dividing it by the incident energy, $e$, viz. $E_{i}=v_{i}/e$. Notice that $E_{i}\in\left[0,1\right]$, where the left and right bounds correspond to when the voxel energy is zero and equal to the incident energy, respectively. To remove the strict bounds, we define $u_{i}=\delta+(1-2\delta)E_{i}$, where $\delta=10^{-7}$, to prevent discontinuities during the logit transformation. Specifically, we use the transformation $x_{i}=\ln u_{i}/(1-u_{i})-\ln\delta/(1-\delta)$, where the second term preserves the zero values in the transformed variable, i.e., when the voxel energy is zero, $v_{i}=0$, the transformed variable $x_{i}=0$.
@@
Besides the reconstruction of the event, the encoded data is also used to train the 4-partite RBM via the Kullback-Liebler (KL) divergence. As in traditional VAEs (full derivations can be found in Ref. [35] and in Appendix A), we optimize the evidence lower bound (ELBO), $\mathcal{L}_{\phi,\theta}(\bm{x})$, to train the Calo4pQVAE. Explicitly, the ELBO function is:
@@
The first term in Eq. (2) represents the reconstruction accuracy and the second term corresponds to the KL divergence. To define a functional form for $p_{\theta}(\hat{\bm{x}}|\bm{z})$, certain assumptions about the distribution are typically made. A common assumption is that the likelihood of the reconstruction is Gaussian distributed, which simplifies to optimizing the mean squared error (MSE).
@@
Maximizing the log-likelihood of $p_{\theta}(\bm{\chi},\bm{\xi}|\bm{z})$ is rather difficult due to the factor containing the Dirac delta function. It may well be possible to replace the Dirac delta function with a smooth parameterized function. Yet, there is a simpler way where we instead mask $\bm{\chi}$ with $\bm{\xi}$, in which case we can neglect the term containing the Dirac delta function. The key point to stress here is that by means of the mask we split the tasks between generating zeroes and non-zeroes, and generating the voxel energy via $\bm{\xi}$ and $\bm{\chi}$, respectively. Therefore, the joint distribution in Eq. (3) becomes  [^2] We removed the $\xi$ coefficient due to normalization.:
@@
Let us focus on the first term on the r.h.s. of Eq. (7). When taking the gradient of the entropy with respect to the model parameters one faces two issues: i) due to the discrete nature of the latent variables $z$, the gradient becomes singular; and ii) given a function evaluated on a random variable, it is ill-defined taking the gradient of the function with respect to the parameters of the random variable’s probability density function. Moreover, taking the gradient of a discrete estimator, $\sum_{\bm{z}\sim q_{\phi}(\bm{z}|\bm{x})}(...)$, with respect to $\phi$ is not well-defined, since the gradient variables appear in the summation argument. The first issue can be easily addressed by simply considering a continuous step-like function, such as a sigmoid, $\sigma(\bullet)$. Specifically, we replace $\bm{z}$ with $\bm{\zeta}$, where $\zeta_{i}=\sigma(\left[\text{encoded data point}\right]_{i}\cdot\beta)$, and $\beta$ is an annealing parameter, such that $\lim_{\beta\rightarrow\infty}\bm{\zeta}=\bm{z}$. The second issue can be addressed by means of the so-called Gumbel trick [37]. The Gumbel trick has become an umbrella term which refers to a set of methods to sample from discrete probabilities or to estimate its partition function. In our case, we simply generate latent variables $\bm{\zeta}$ via
@@
The first term in the r.h.s. is the energy function averaged over the approximate posterior, whereas the second term is the free energy, which is independent of the encoded samples. The free energy can be replaced by the internal energy, $U$, minus the entropy, $S$, times the temperature, viz. $U-TS$. We use the high temperature gradient approximation to replace the second term in the r.h.s. of Eq. (10) with the internal energy, which is a common practice in approximating the free energy when training RBMs. In the high temperature gradient approximation, as the energy in the system saturates, the specific heat converges to zero. We demostrate in Appendix D that when $\langle E(\bm{z})\rangle\langle\partial E(\bm{z})/\partial\phi\rangle=\langle E(\bm{z})\partial E(\bm{z})/\partial\phi\rangle$, the specific heat is zero, which allows us to rewrite the previous Equation as
@@
We then propose a Gibbs sampling process to sample from the joint distribution $p(\mathbf{v},\mathbf{h},\mathbf{s},\mathbf{t})$ of the 4-partite RBM. Notice that the joint distribution can be factorized as $p(\mathbf{v},\mathbf{h},\mathbf{s},\mathbf{t})=p(\mathbf{v}|\mathbf{h},\mathbf{s},\mathbf{t})p(\mathbf{h},\mathbf{s},\mathbf{t})$. Extending the approach used in 2-partite RBM, we perform Markov chain Monte Carlo by sequentially updating each partition conditioned on the remaining, i.e., the prior in step $n+1$ is deemed the posterior in step $n$. The Gibbs sampling process, at iteration $n$, is done in the following steps:
@@
Notice that one block Gibbs sampling step corresponds to four sampling steps. By repeating this process iteratively, we generate samples that approximate the joint distribution.
We have shown how to generalize an RBM to a four-partite graph. It is important to emphasize that, as with any deep generative model, the features of the dataset are expressed in latent space. In this context, each state in latent space encodes specific features of the dataset. Different techniques such as t-SNE [40] and LSD [41], can illustrate how dataset features are embedded in latent space after training. These methods allow for the conditioning of the prior on specific features. There is a rather straightforward way to condition the prior on given dataset parameters during training, which is widely used across different frameworks. Generally, this involves using specific channels to condition the prior and specifics depend on the actual framework. An accurate and robust method to sample from latent space constrained to specific features is crucial for practical applications in generative deep models. In the present context, a calorimeter surrogate requires input related to particle type, incident particle energy and incident angle, among other parameters. In the following, we show how to condition the 4-partite RBM sampling process.
@@ II.3.1 Conditioned 4-partite Restricted Boltzmann Machine
The block Gibbs sampling procedure is the same as before via Eqs. (17) although we skip Eq. (17a). The authors in [42] caution on the use of contrastive divergence when training a conditioned RBM due to the possibility of vanishing gradient even when the conditioned RBM might not be fully trained. This happens when mixing times are larger than the number of Gibbs sampling steps. Despite the previous, we use contrastive divergence in our framework and we show how the estimated log-likelihood saturates after $250$ epochs. We further elaborate and discuss this approach in the discussion section.
@@ II.4 Quantum Annealers
A quantum annealer (QA) is an array of superconducting flux quantum bits with programmable spin–spin couplings and biases with an annealing parameter [43]. The motivation for QAs comes from the adiabatic approximation [44], which asserts that if a quantum system is in an eigenstate of its Hamiltonian (which describes the total energy of the system), and the Hamiltonian changes slowly enough, then the system will remain in an eigenstate of the Hamiltonian, although the state itself may change (see Appendix F for a formal derivation). QAs were initially thought of as a faster method to find the ground state of complex problems that could be mapped onto a Hamiltonian $H$ [45]. This can be done by initializing the system in the ground state of some Hamiltonian $H_{0}$, which is easy to prepare both theoretically and experimentally. In addition, by design the commutator $[H,H_{0}]\neq 0$. The QA interpolates between the two Hamiltonians via
@@
Notice that the RBM data are binary vectors such that each element can take value of 00 or $1$, whereas the qubits can have values $-1$ or $1$. Hence, to map the RBM onto the QA one redefines the RBM variables $x_{i}\rightarrow(\sigma_{z}^{(i)}+1)/2$ and rearrange the terms in the RBM Hamiltonian to be mapped onto the QA (this is explicitly shown in Appendix F). The previous variable change will lead to an energy offset between the RBM and the QA, which henceforth we neglect. It is important to stress that the QA is in a heat bath with temperature $T_{QA}\lesssim 15mK$ [53], while the annealing parameter has an upper bound $\mathcal{B}(1)\approx 5.0\cdot 10^{-24}J$ [54], which implies that the prefactor $\beta_{QA}=\frac{\mathcal{B}(1)}{2k_{B}T_{QA}}\approx 12$ while in the RBM by design $\beta_{RBM}=1$. To ensure that both, the RBM and the QA describe the same Boltzmann distribution with the same temperature, one can either re-scale the QA Hamiltonian by $1/\beta_{QA}$ or the RBM Hamiltonian by $\beta_{QA}$. Currently, there is not a way to measure the QA prefactor, however there are different ways to estimate it [47, 55]. In addition, there is not a direct way to control this prefactor. The most common way to estimate this prefactor is via the Kullback-Liebler divergence between the QA and the RBM distributions, whereby one introduces a tuning parameter $\beta$ as a prefactor in the RBM distribution which is tuned to minimize the Kullback-Liebler divergence. It is in the process of estimating the $\beta_{QA}$ that we also indirectly control the parameter by rescaling the Hamiltonian. The previous can be mathematically expressed as
@@
The previous converges when the average energy from the RBM matches that from the QA, in which case, $\beta=\beta_{QA}$. Due to coupling-dependent temperature fluctuations, $\beta_{QA}$ is expected to change in the process of training the model. For this reason the method described by Eq. (22) should be employed, in principle, after each parameter update. Notice that optimizing $\beta$ in Eq. (22) requires samples from the RBM, furthermore, it is the classical RBM temperature which is being tuned to match the distribution of the QA. In our case, we want the opposite, to fit the QA distribution to that of the RBM and not the other way around. To address this point, $H(x)$ is replaced with an annealed Hamiltonian $H(x,\beta)=H(x)/\beta$, i.e., instead of tuning the temperature in the classical RBM, to match the distribution in the QA, one iteratively rescales the Hamiltonian in the QA to effectively tune the QA’s temperature to match that of the classical RBM. This approach, as already mentioned, is well-known [47, 55, 31] and has been proven to be empirically robust yet slow to converge. We therefore propose a different protocol which empirically converges faster than the KL divergence:
@@
Up to this point, we have detailed the process of replacing the RBM with the QA in the Calo4pQVAE. As previously mentioned, we utilize a conditioned RBM to enable the sampling of showers with specific characteristics. In the next section, we will explore various approaches for conditioning the QA, aiming to achieve targeted sampling of particle showers with desired features.